What are [H3O+], [OH], and pH in a solution with a pOH of 3.44?

[H3O+] = ___× 10M

[OH] = ____× 10M

pH = ____
Answer #1

Ans :

pH + pOH = 14

pH = 14 - pOH

= 14 - 3.44

= 10.56

[H3O+] = 10-pH

= 10-10.56

= 2.75 x 10-11 M

[OH-] = 10-pOH

= 10-3.44

= 3.63 x 10-4 M

pH = 10.56